lauren is 4 years older than meghan. alyssa is 4 years younger than meghan. the sum of their ages is 42. how old is each girl?

Found 2 solutions by ankor@dixie-net.com, greenestamps:
Answer by ankor@dixie-net.com(22740)

lauren is 4 years older than meghan.
L = m + 4
alyssa is 4 years younger than meghan.
a = m - 4
the sum of their ages is 42.
L + m + a = 42
Replace L with (m+4); replace a with (m-4)
(m+4) + m + (m-4) = 42
3m = 42
m = 42/3
m = 14 yrs is Meghan's age
then
14 + 4 = 18 yrs is Lauren's age
and
14 - 4 = 10 yrs is Alyssa's

Answer by greenestamps(13203)



Lauren is 4 years older than Meghan; Alyssa is 4 years younger than Meghan.

That means the three ages form an arithmetic progression (constant difference of 4 between the ages).

That means the middle age is the average of the three ages; 42/3 = 14.

ANSWER: Meghan (middle) is 14; Lauren is 4 years older, 18; Alyssa is 4 years younger, 10.